OpenStudy (anonymous):

Guys I'm sad because I haven't done my math in like 6 months because I hate radicals. :( Help me pwease? It's probably super easy but my lesson doesn't explain much and my teacher refuses to help. Change this radical to an algebraic expression with fractional exponents. 3√a

OpenStudy (unklerhaukus):

\[\sqrt x=x^{1/2}\]

OpenStudy (anonymous):

a^(1/3). Basically the rule here is fairly simple. Say you have 2√x^3 (square root of x to the power of 3. What you do is take the "to the power of" portion and divide it by what kind of root it is (ie. squared is 2, cubed is 3...etc). So what you get is x^3/2

OpenStudy (anonymous):

What about the 3? I'm really sorry I just don't know and it's really stressing me out :( I'm lame.

OpenStudy (unklerhaukus):

\[\sqrt[n]x=x^{1/n}\]

OpenStudy (anonymous):

Ohhhhhh Thanks @Evgeny I'm in Alg II and I have an A but I almost failed Alg I and my teacher was like "Well you should've learned that in Alg I...so figure it out" :/

OpenStudy (unklerhaukus):

\[\sqrt[n]{x^m}=(\sqrt[n]x)^m=x^\frac mn\]

OpenStudy (anonymous):

Thank you @UnkleRhaukus I will do my best to remember all of this. I will probably still ask for more help haha

OpenStudy (unklerhaukus):

thats why we are here. \[\color{teal}{\ddot\smile}\]

OpenStudy (anonymous):

Would √b just be b^2? Or something...

OpenStudy (hba):

@yololol No you are wrong \[\sqrt{b}=(b)^{1/2}\]

OpenStudy (anonymous):

It doesn't give me a fraction as an option, idt...hold on

OpenStudy (anonymous):

Okay, it doesn't give it as an option to be an exponent so maybe they'll just assume it's an exponent and I'll hope for the best. Thank you! Sorry about my dumb questions. OMG it's Joker!! :D xD

OpenStudy (precal):

not sure which one you meant, so I wrot both versions out for you